Greg spent 1/4 hour (h) responding to emails and 5/8 h returning phone calls. How much time did he spend on the two activities?
Answer by rajagopalan(174)

Greg spent 1/4 hour (h) responding to emails
5/8 h returning phone calls.
on the two activities he spent 1/4 plus 5/8 h
=(1/4)+5/8)
=(4/16)+(10/16) h
=14/16 h
=7/8 h
okay